OpenStudy (anonymous):

multiply and simplify 285 then in parentasees its a fraction of 1 over -9

jimthompson5910 (jim_thompson5910):

\[\large 285\left(\frac{1}{-9}\right)\] \[\large \frac{285}{-9}\] \[\large \frac{95\times3}{-3\times3}\] \[\large \frac{95\times\cancel{3}}{-3\times\cancel{3}}\] \[\large \frac{95}{-3}\] \[\large -\frac{95}{3}\] So \[\large 285\left(\frac{1}{-9}\right)=-\frac{95}{3}\]